OpenStudy (anonymous):

3x+2y=23 x+5z= -7 2y-z+16 How do you do this??? PLEASE HELP

OpenStudy (bahrom7893):

Some substitution should do the trick: From equation 2: x = -7-5z

OpenStudy (anonymous):

i know you use substitution but i still get stuck T.T

OpenStudy (bahrom7893):

Plugging that instead of x into equation 1 gives us: 3(-7-5z) + 2y = 23, a bit of simplification will give us: -21-15z+2y=23 -15z+2y=23+21 -15z+2y=44

OpenStudy (bahrom7893):

By the way is the third equation this: 2y-z=16 (= instead of +)

OpenStudy (bahrom7893):

Hello?

OpenStudy (anonymous):

oh srry let me try it out hold on

OpenStudy (anonymous):

oh no it is +

OpenStudy (anonymous):

hello???

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

It must have been a typo, that's impossible to solve otherwise.

OpenStudy (anonymous):

so what do i do after i plugged in the x?

OpenStudy (anonymous):

ohh no no yes yes it is = i was looking at the wrong + SORRY

OpenStudy (anonymous):

so what do i do now? its narrowed down to -15z+2y=44 now what?

OpenStudy (anonymous):

@bahrom7893

OpenStudy (bahrom7893):

Plugging that instead of x into equation 1 gives us: 3(-7-5z) + 2y = 23, a bit of simplification will give us: -21-15z+2y=23 -15z+2y=23+21 -15z+2y=44 From the third equation find z: 2y-z=16 2y-16=z z=2y-16 Plugging this into the equation above: -15(2y-16)+2y=44

OpenStudy (bahrom7893):

Solve for y, then find z, then find x

OpenStudy (bahrom7893):

z = 2y-16 x = -7-5z

OpenStudy (anonymous):

how do i solve for y though? i f i keep having two variables???

OpenStudy (anonymous):

OHH wait wait nvm i get it i get it haha okay let me solve the whole thing and see if i get an answer :) THANK YOU SO MUCH :)

OpenStudy (bahrom7893):

Go for it :)

OpenStudy (anonymous):

I got the answer! It's 8 :) and it was one of the choices! Thank you! :) Do you mind helping me again later??? :)

OpenStudy (anonymous):

@bahrom7893